OpenStudy (anonymous):

Find the center and radius of the circle. 3x^2+3y^2-12x-24+12=0 Could u please explain in steps

OpenStudy (anonymous):

\( \huge 3x^2+3y^2-12x-\underbrace{24}_{24y}+12=0 \) is this what you meant.. ? :)

OpenStudy (anonymous):

@tylerjh still here?

OpenStudy (anonymous):

first of all, you need to know its formula. \( \LARGE (x-h)^2+(y-k)^2=r^2\) where Center \(\large h,k\;\;\;\)and radius \(r\)

OpenStudy (anonymous):

\[\LARGE 3x^2+3y^2-12x-24y+12=0\] \[\LARGE (3x^2-12x)+(3y^2-24y)=-12\] \[\LARGE 3(x^2-4x)+3(y^2-8y)=-12\] \[\LARGE \frac{3(x^2-4x)}{3}+\frac{3(y^2-8y)}{3}=-\frac{12}{3 }\] \[\LARGE \frac{\cancel{3}(x^2-4x)}{\cancel3}+\frac{\cancel 3(y^2-8y)}{\cancel 3}=-4\] \[\LARGE (x^2-4x)+(y^2-8y)=-4\] what do we need to do here to complete the square.. ? :)

OpenStudy (anonymous):

@tylerjh I'm not going to do your homework for you... if you don't collaborate, I'll leave !

OpenStudy (anonymous):

im very sorry i left the computer to help my mother with moving. Sorry man. I really am. That wont happen again

OpenStudy (anonymous):

Ohh never mind. I thought you were sitting there watching me do it. Ok To complete the square you have to do this:  so you need to add 4 and -4 to x's part (so the value won't change.. 4-4=0 ) so we have: \[\LARGE (x^2-4x+4)-4+(y^2-8y)=-4\] now we'll do the same with y's \(\LARGE \frac{-8}{2}=-4\longrightarrow (-4)^2=16\) so we'll add 16 and -16 \[\Large (x^2-4x+4)-4+(y^2-8y+16)-16=-4\] now we move -16 and -4 to the right side... \[\Large (x^2-4x+4)+(y^2-8y+16)=-4+16+4 \] and we sum\simplify them... \[\Large (x^2-4x+4)+(y^2-8y+16)=\cancel{-4}+16\cancel{+4} \]